How is the cost of a lump-sum purchase allocated to the individual assets acquired? Provide a...

  1. How is the cost of a lump-sum purchase allocated to the individual assets acquired? Provide a simple example of how the allocation would be performed if your company paid $1,000,000 cash in total for a restaurant (land, building, and equipment), and the appraised value was actually $1,200,000 for the 3 asset components $400,000, $500,000, and $300,000, respectively.
  2. Does the balance in the Accumulated Depreciation – Machinery account represent funds to replace the machinery when it wears out? If not, what does it represent?
  3. Assume you work for a company in sales and therefore are entitled to not only hourly pay, but commission and bonus as well. You work 40 hours in a payroll week. Calculate your gross pay for your company if you make $x per hour (you select your rate!), earn a 5% bonus (of gross pay), and earn 10% commission (of gross pay).
  4. Unfortunately, you do not get to keep all your Gross Pay. Assuming you have the following withholding deductions taken out of your pay, now calculate your Net Pay.
Withholding deductions:
Employee income tax (20%)
Employee OASDI tax (see book for how to calculate)
Employee Medicare tax (1.45%)
Employee health insurance 75
Employee contribution to United Way 20
Total withholdings:
Answer #1

The cost of a lump-sum purchase will be allocated to the individual assets acquired on the basis of their appraised value. For example, if the company paid $1,000,000 cash in total for a restaurant (land, building, and equipment), and the appraised value was actually $1,200,000 for the 3 asset components $400,000, $500,000, and $300,000, respectively, then this cost of $1,000,000 will be allocated on individual asset as follows:

Asset Land Building Equipment
Appraised value 400000 500000 300000
Purchase Value =400000/120000*100000 =500000/120000*100000 =300000/120000*100000
=333333.333333333 =416666.666666667 =250000

No, the balance in the Accumulated Depreciation – Machinery account does not represent funds to replace the machinery when it wears out. Instead, it represents the amount of decline in the asset.

Calculation of Net Pay $
Gross Pay @ 100 per hour =40*100 $4,000.00
Add: 5% bonus =4000*5/100 $200.00
Add: 10% commission =4000*10/100 $400.00
Total Pay (A) $4,600.00
Withholding deductions:
   Employee Income Tax (20%) =4600*20/100 $920.00
   Employee OASDI Tax =4600*6.2/100 $285.20
   Employee Medicare Tax (1.45%) =4600*1.45/100 $66.70
   Employee Health Insurance $75.00
   Employee Contribution to United Pay $20.00
Total Withholdings (B) $1,366.90
Net Pay (A-B) $3,233.10